RC Meeting and Runplan, Thursday, Sept 20

From HallCWiki
Jump to: navigation, search
  • Status
    • Target cool down had some issues due to cryo. Dave Meekins is down in the Hall and working with cryo to resolve.
    • Target is setup so that selection of liquid targets is disabled.
    • Hall lock up sometime this afternoon once target is set. All checklists are complete. Hall is otherwise ready. Old HMS rotation motor replaced today and tested 35-15deg, SHMS has new motor still installed. SHMS pointing survey completed.
    • Possible beam in hall tonight. N. Linac fault is going to cause delay of at least 4 hours. PD is going to update me later this evening on the shift tomorrow. Day shift takers should plan to show up unless otherwise notified (may not be much action in the first part of the day).
    • SHMS is parked at 6.5 degrees, HMS at 15.01 degrees.
    • Will request an access Monday morning for Dave to condense target and survey.
  • Runplan
    • Starting at 5 pass. Updated plan here: [1]